Suppose that after 10 years of service, 30% of computers have problems with motherboards (MB), 20% have problems with hard drives (HD), and 15% have problems with both MB and HD. What is the probability that a 10-year old computer still has fully functioning MB and HD? (10)
Answer by ikleyn(52905)

.

    P(MB or HD) = P(MB) + P(HD) - P(both) = 0.3 + 0.2 - 0.15 = 0.35.


    P(a 10-years old computer is OK) = 1 - 0.35 = 0.65 = 65%.    ANSWER

Solved.